Two numbers total 66 and have a difference of
32. Find the two numbers.

Answer by addingup(3677)

x + y = 66
x - y = 32 add y on both sides and you get:
x = 32 + y now substitute for x in the first equation:
x + y = 66
32 + y + y = 66
2y = 34
y = 17
and
x - y = 32
x - 17 = 32
x = 49
----------------------
Check:
x + y = 66
49 + 17 = 66 Correct